OpenStudy (anonymous):

When 1600 is increased by 20% and the resulting number is decreased by y% the answer is 1536. Find the value of y?

OpenStudy (anonymous):

1600 increased by 20% is 1920.

OpenStudy (anonymous):

Now 1920 x (y/100) = 1536 Am I right?

OpenStudy (anonymous):

Now when i solve for y I get 80 which is wrong.

OpenStudy (anonymous):

So how do I solve

OpenStudy (campbell_st):

oops reduction is 384 so 384/1920 * 100 = 20% y = 20% reduction

OpenStudy (anonymous):

1920 - 1920*(y/100) = 1536

OpenStudy (anonymous):

Thank you @medisynergi and @campbell_st

OpenStudy (anonymous):

Solve the following for y: (1600*1.2)*(1 - y) = 1536 y = 0.2 or 20% off